# Low-Altitude ISAC with Rotatable Active and Passive Arrays

### v1

@@ III-D1 Gradient computation
Let $\mathcal{D}\triangleq\mathcal{D}_{\text{B}}\times\mathcal{D}_{\text{R}}$ be the box of angle limits in (P4) with
$\mathcal{D}_{\text{A}}=\{\boldsymbol{r}^{\text{A}}:\boldsymbol{r}_{\text{lb}}^{\text{A}}\leq\boldsymbol{r}^{\text{A}}\leq\boldsymbol{r}_{\text{ub}}^{\text{A}}\},$ (82)
and denote $\boldsymbol{r}_{\text{lb}}\!\triangleq\![(\boldsymbol{r}_{\text{lb}}^{\text{B}})^{T}\!,(\boldsymbol{r}_{\text{lb}}^{\text{R}})^{T}]^{T}$, and $\boldsymbol{r}_{\text{ub}}\!\triangleq\![(\boldsymbol{r}_{\text{ub}}^{\text{B}})^{T}\!,(\boldsymbol{r}_{\text{ub}}^{\text{R}})^{T}]^{T}$. Let $\Pi_{\mathcal{D}}(\cdot)$ be the Euclidean projection onto $\mathcal{D}$. We perform projected gradient ascent on $F$ with the projected step
$\boldsymbol{\beta}\left(\alpha\right)=\Pi_{\mathcal{D}}\left(\boldsymbol{r}+\alpha\nabla_{\boldsymbol{r}}F\left(\boldsymbol{r}\right)\right)-\boldsymbol{r},$ (83)
for any step size $\alpha>0$. Define the projected-gradient mapping $\boldsymbol{H}_{\alpha}\left(\boldsymbol{r}\right)\triangleq\frac{1}{\alpha}\boldsymbol{\beta}\left(\alpha\right)$, for any $\alpha>0$, $\boldsymbol{r}^{\star}$ is a first-order stationary point of (P4) if and only if $\boldsymbol{H}_{\alpha}\left(\boldsymbol{r}^{\star}\right)=\boldsymbol{0}$, according to the KKT condition. Hence $\|\boldsymbol{H}_{\alpha}(\boldsymbol{r})\|_{2}$ is a valid criticality measure for stopping. Because $\mathcal{D}$ is a box, the projection (83) reduces to the componentwise clipping
$\boldsymbol{\beta}^{(\ell)}(\alpha)=\text{clip}(\boldsymbol{r}^{\left(\ell\right)}+\alpha\nabla_{\boldsymbol{r}}F^{(\ell)}(\boldsymbol{r}))_{[r_{\text{lb}},r_{\text{ub}}]}-\boldsymbol{r}^{(\ell)},$ (84)
where the clip $\text{clip}\left(x\right)_{\left[y,z\right]}=\min\left\{\max\left\{x,y\right\},z\right\}$ acts elementwise. Given the previous $\left(\boldsymbol{r},F\left(\boldsymbol{r}\right),\nabla F\left(\boldsymbol{r}\right)\right)^{(\ell-1)}$, the stepsize uses Barzilai–Borwein initialization
@@
$F\left(\boldsymbol{r}+\boldsymbol{\beta}\left(\alpha\right)\right)\geq F\left(\boldsymbol{r}\right)+c_{1}\nabla_{\boldsymbol{r}}F\left(\boldsymbol{r}\right)^{T}\boldsymbol{\beta}\left(\alpha\right),$ (86)
with $c_{1}\!\in\!(0,1)$ and the backtracking using $\alpha\!\leftarrow\!c_{2}\alpha$, $c_{2}\in(0,1)$. The iteration terminates when it meets the stopping criteria that the projected gradient norm is small, i.e., $\|\boldsymbol{H}_{\alpha}(\boldsymbol{r}^{(\ell)})\|_{2}\ \leq\ \varepsilon_{\text{pg}}$, or the relative step is small, i.e., $\frac{\|\boldsymbol{r}^{(\ell)}-\boldsymbol{r}^{(\ell-1)}\|_{2}}{\max\{1,\|\boldsymbol{r}^{(\ell-1)}\|_{2}\}}\ \leq\ \varepsilon_{\text{rel}}$. For box constraints, $\boldsymbol{H}_{\alpha}(\boldsymbol{r}^{(\ell)})$ admits the following componentwise form for $\imath$-th entry independent of $\alpha$
$\!\!\!\big[\boldsymbol{H}_{\alpha}(\boldsymbol{r}^{(\ell)})\big]_{\imath}\!=\!\!\begin{cases}[\nabla_{\boldsymbol{r}}F^{\left(\ell\right)}(\boldsymbol{r})]_{\imath},\!&\!\!\!r_{\text{lb},i}\!\!<r_{i}\!<r_{\text{ub},i}\!,\\ \max\{0,[\nabla_{\boldsymbol{r}}F^{(\ell)}(\boldsymbol{r})]_{\imath}\},\!&\!\!\!r_{\imath}=r_{\text{lb},\imath},\\ \min\{0,[\nabla_{\boldsymbol{r}}F^{(\ell)}(\boldsymbol{r})]_{\imath}\},\!&\!\!\!r_{\imath}=r_{\text{ub},\imath}.\end{cases}\!\!\!\!$ (87)
The gradient cost is dominated by Jacobians (80) of rotation-related steering vectors and derivatives (78) with matrix-chain products. Because the directional-gain model is piecewise smooth, we use subgradients at the visibility boundary and a smooth gating for stable line search.
@@
1:  Input: fixed $(\boldsymbol{W},\boldsymbol{\theta})$, box $\mathcal{D}$, tolerances $\varepsilon_{\mathrm{pg}},\varepsilon_{\mathrm{rel}}>0$, and maximum iterations $I_{\text{PGD}}^{\text{max}}$.
2:  Initialization: initial point $\boldsymbol{r}^{(0)}\in\mathcal{D}$
@@
13:   Compute projected gradient mapping $\boldsymbol{H}_{\alpha}(\boldsymbol{r}^{(\ell)})$ via (87).
14:   if $\|\boldsymbol{H}_{\alpha}(\boldsymbol{r}^{(\ell+1)})\|_{2}\leq\varepsilon_{\mathrm{pg}}$ or $\dfrac{\|\boldsymbol{r}^{(\ell+1)}-\boldsymbol{r}^{(\ell)}\|_{2}}{\max\{1,\|\boldsymbol{r}^{(\ell)}\|_{2}\}}\leq\varepsilon_{\mathrm{rel}}$ then
@@
18:  Output: $\boldsymbol{r}^{\mathrm{B}}\leftarrow\boldsymbol{r}^{(\ell+1)}_{1:3}$, $\boldsymbol{r}^{\mathrm{R}}\leftarrow\boldsymbol{r}^{(\ell+1)}_{4:6}$.
III-D3 Overall algorithm under penalty-based AO framework for original problem 1
@@
Algorithm 4 summarizes the overall projected gradient method. For the per-iteration computational complexity, directly forming $\boldsymbol{B}$ and the Jacobians as dense $M\times N$ matrices yields $\mathcal{O}(PMN)$ computations per rotation axis, and evaluating the sensing gradients over $A$ grids would add $\mathcal{O}(AMN)$. However, using the rank-$P$ factorization for $\boldsymbol{B}$ and its derivatives, all costly matrix–vector products, involving $\boldsymbol{B}$ and $\partial\boldsymbol{B}/\partial r_{\imath}^{\mathrm{A}}$, are computed without materializing $M\times N$ blocks and costing $\mathcal{O}(NP)$ plus $\mathcal{O}(MP)$. With these reuses, one PGD iteration with full gradient of $F$ has the leading cost: i) precomputation on steering vectors and their Jacobians with $\mathcal{O}\big(P(M+N)\big)$, ii) sensing and communication channel–Jacobian products $\mathcal{O}\big((A+K)P(N+M)\big)$, iii) use $\boldsymbol{W}(\boldsymbol{W}^{H}\boldsymbol{f})$ to apply $\boldsymbol{W}\boldsymbol{W}^{H}$ in (71) and (81) with $\mathcal{O}\!\big((M+K)M(A+K)\big)$. Therefore, the per-iteration complexity scales as $\mathcal{O}\big(P(M+N)(A+K+1)+(M+K)M(A+K)\big)$ with a constant coefficient affected by backtracking searches.
III-E Overall algorithm
Finally, we adopt a penalty-assisted AO-based algorithm framework over $(\boldsymbol{W},\boldsymbol{\theta},\boldsymbol{r}^{\mathrm{B}},\boldsymbol{r}^{\mathrm{R}})$, resulting in an overall algorithm with nested loops summarized in Algorithm 4, where we obtain a suboptimal solution for the original problem (P1). At the $\delta$-th outer iteration, with the other blocks fixed, we first solve (P2) by Algorithm 1 to update $\boldsymbol{W}^{(\delta+1)}$; next solve (P3) by Algorithm 2 to update $\boldsymbol{\theta}^{(\delta+1)}$; then solve (P4) by Algorithm 3 (projected-gradient over the box $\mathcal{D}$) to obtain $(\boldsymbol{r}^{\mathrm{B}},\boldsymbol{r}^{\mathrm{R}})^{(\delta+1)}$. After the three inner loops, the penalty weight is refreshed as $\rho^{(\delta+1)}\leftarrow\gamma_{\rho}\rho^{(\delta)}$ if the MSE constraint is violated (i.e., $\mathrm{MSE}>\eta_{\text{th}}$); otherwise $\rho^{(\delta+1)}=\rho^{(\delta)}$. Each inner solver uses its own line search/tolerances to guarantee ascent of the penalized objective and is warm-started from the previous iterate. The outer loop terminates when the fractional increase of the penalized objective in (P1.1) falls below $\varepsilon_{\mathrm{AO}}$ and the MSE constraint is satisfied ($\mathrm{MSE}\leq\eta_{\text{th}}$).
Algorithm 4 Penalty-assisted AO-based Algorithm for (P1)
1:  Input: $\boldsymbol{u}_{k,l}^{\text{BU}},\boldsymbol{u}_{p}^{\text{RB}},\boldsymbol{u}_{p}^{\text{BR}},\boldsymbol{u}_{k,q}^{\text{RU}},\boldsymbol{u}_{a}^{\text{BT}},\boldsymbol{u}_{a}^{\text{RT}}$, $\gamma_{\rho}\!>\!0$, and $\mathcal{P}_{\text{d}}\left(\vartheta_{a},\varphi_{a}\right)$ .
2:  Initialization: $\boldsymbol{W}^{\left(0\right)},\boldsymbol{\theta}^{\left(0\right)},\left(\boldsymbol{r}^{\text{A}},\boldsymbol{r}^{\text{B}}\right)^{\left(0\right)}$, $\rho^{(0)}$ and $\delta=0$
@@ III-E Overall algorithm
4:   Solve subproblem (P2) via Algorithm 1 to obtain $\boldsymbol{W}^{\left(\delta+1\right)}$.
5:   Solve subproblem (P3) via Algorithm 2 to obtain $\boldsymbol{\theta}^{\left(\delta+1\right)}$.
6:   Solve subproblem (P4) via Algorithm 3 to obtain $(\boldsymbol{r}^{\text{A}}\!,\!\boldsymbol{r}^{\text{B}})^{(\delta\!+\!1)}$.
7:   if Current MSE value violates the threshold $\eta_{\text{th}}$ then
8:    $\rho^{(\delta+1)}=\gamma\rho^{(\delta)}$.
9:   else
10:    $\rho^{(\delta+1)}=\rho^{(\delta)}$.
11:   end if
12:   $\delta\leftarrow\delta+1$.
13:  until the fractional increase of the objective value of problem (P1.1) is smaller than a threshold $\varepsilon_{\text{AO}}$ and the MSE is below the threshold $\eta_{\text{th}}$.
14:  Output: $\boldsymbol{W}\leftarrow\boldsymbol{W}^{(\delta)},\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}^{(\delta)},\boldsymbol{r}^{\text{A}}\leftarrow\boldsymbol{r}^{\text{A},(\delta)}$, $\boldsymbol{r}^{\text{B}}\leftarrow\boldsymbol{r}^{\text{B},(\delta)}$.
@@ IV Simulation Results
This section presents numerical results for the proposed rotatable BS–RIS ISAC design. Unless otherwise stated, the BS is equipped with a $M=M_{x}M_{y}=2\times 2$ antennas by default with inter-element spacing $d_{x}=d_{y}=\lambda/2$ for a carrier wavelength $\lambda=0.1$ m. The RIS is modeled as a $6\times 6$ UPA with spacing $e_{x}=e_{y}=\lambda/2$, located at $(10,0,0)$ m relative to the BS. The unrotated BS and RIS boresights are aligned with the positive $z$-axis, and their orientations are parameterized by rotation vectors $\boldsymbol{r}^{\mathrm{B}}$ and $\boldsymbol{r}^{\mathrm{R}}$. The small-scale fading of all links follows a geometry-based multipath model: each BS–user and RIS–user channel comprises $L_{k}=Q_{k}=2$ paths with independently generated azimuth and elevation angles, uniformly distributed over $\pi,\pi]$ and $\pi/3,\pi/3]$, respectively, and complex Gaussian path gains with normalized variance. The cascaded BS–RIS channel contains $P=2$ paths with angles and gains drawn independently. Thermal noise at each user is modeled as circularly symmetric complex Gaussian with variance $\sigma_{k}^{2}=10^{-6}$. For sensing, the BS forms a monostatic beampattern over a two-dimensional angular grid with $A_{\mathrm{az}}=11$ azimuth and $A_{\mathrm{el}}=6$ elevation samples, uniformly spanning $\phi\in[-\pi,\pi]$ and $\vartheta\in[-\pi/4,\pi/4]$. The desired sensing pattern corresponds to two spotlight-like angular sectors, and the target threshold of sensing MSE is $\eta_{\mathrm{th}}=0.2$. The total BS transmit power is $P_{\mathrm{B}}=30$ dBm.
We adopt both half-space isotropic and directional element patterns. In the directional case, the BS and RIS element gains follow the piecewise model (11), while $G_{0,\mathrm{B}}=G_{0,\mathrm{R}}=2$ are constant in the isotopic case. For each random channel realization, the proposed alternating optimization algorithm jointly updates the BS precoder $\boldsymbol{W}$, the RIS phase vector $\boldsymbol{\theta}$, and the rotations $(\boldsymbol{r}^{\mathrm{B}},\boldsymbol{r}^{\mathrm{R}})$ so as to maximize the downlink sum rate $\sum_{k=1}^{K}\log_{2}(1+\mathrm{SINR}_{k})$ under the sensing MSE constraint. To assess the gain of spatial geometry control and RIS assistance, we compare the fully rotatable BS–RIS architecture (“Rot-BS+Rot-RIS”) with four benchmarks: (i) only the BS is rotatable while the RIS orientation is fixed (“Rot-BS+Fix-RIS”); (ii) only the RIS is rotatable (“Fix-BS+Rot-RIS”); (iii) both BS and RIS are fixed (“Fix-BS+Fix-RIS”); and two non-RIS baseline (iv) (“Rot-BS+No-RIS”) and (v) (“Fix-BS+No-RIS”), where the RIS–user link is disabled by setting the RIS–user path gains to zero. All reported curves are obtained by averaging the communication and sensing metrics over a large number of independent channel realizations.

### v2

@@ III-D1 Gradient computation
Let $\mathcal{B}$ be the box of array rotation angle limits in (P4)
$\mathcal{B}=\{\boldsymbol{r}:\boldsymbol{r}_{\text{lb}}\preceq\boldsymbol{r}\preceq\boldsymbol{r}_{\text{ub}}\},$ (83)
where we denote $\boldsymbol{r}_{\text{lb}}\triangleq[(\boldsymbol{r}_{\text{lb}}^{\text{B}})^{T},(\boldsymbol{r}_{\text{lb}}^{\text{R}})^{T}]^{T}$ and $\boldsymbol{r}_{\text{ub}}\triangleq[(\boldsymbol{r}_{\text{ub}}^{\text{B}})^{T},(\boldsymbol{r}_{\text{ub}}^{\text{R}})^{T}]^{T}$, and here we let $\boldsymbol{r}_{\text{lb}}^{\text{A}}=[r_{\mathrm{x},\min}^{\text{A}},r_{\mathrm{y},\min}^{\text{A}},r_{\mathrm{z},\min}^{\text{A}}]$ and $\boldsymbol{r}_{\text{ub}}^{\text{A}}=[r_{\mathrm{x},\max}^{\text{A}},r_{\mathrm{y},\max}^{\text{A}},r_{\mathrm{z},\max}^{\text{A}}]$ corresponding to (24d) and (24e), respectively, for brevity. Let $\Pi_{\mathcal{B}}(\cdot)$ be the Euclidean projection onto $\mathcal{B}$. We perform projected gradient ascent of $F\left(\boldsymbol{r}\right)$ on the feasible region with the projected step
$\boldsymbol{\beta}\left(\alpha\right)=\Pi_{\mathcal{B}}\left(\boldsymbol{r}+\alpha\nabla_{\boldsymbol{r}}F\left(\boldsymbol{r}\right)\right)-\boldsymbol{r},$ (84)
for any step size $\alpha>0$. Define the projected-gradient mapping $\boldsymbol{\upsilon}_{\alpha}\left(\boldsymbol{r}\right)\triangleq\frac{1}{\alpha}\boldsymbol{\beta}\left(\alpha\right)$, for any $\alpha>0$, $\boldsymbol{r}^{\star}$ is a first-order stationary point of (P4) if and only if $\boldsymbol{\upsilon}_{\alpha}\left(\boldsymbol{r}^{\star}\right)=\boldsymbol{0}$, according to the KKT condition. Hence $\|\boldsymbol{\upsilon}_{\alpha}(\boldsymbol{r})\|_{2}$ is a valid criticality measure for stopping. Because $\mathcal{B}$ is a box, the projection (84) reduces to the component-wise clipping
$\boldsymbol{\beta}^{(\ell)}(\alpha)=\mathrm{clip}(\boldsymbol{r}^{\left(\ell\right)}+\alpha\nabla_{\boldsymbol{r}}F^{(\ell)}(\boldsymbol{r}))_{[r_{\text{lb}},r_{\text{ub}}]}-\boldsymbol{r}^{(\ell)},$ (85)
where the clip operation is defined by $\mathrm{clip}\left(\mathfrak{a}\right)_{\left[\mathfrak{b},\mathfrak{c}\right]}=\min\left\{\max\left\{\mathfrak{a},\mathfrak{b}\right\},\mathfrak{c}\right\}$ that acts element-wise. Given the previous $\left(\boldsymbol{r},F\left(\boldsymbol{r}\right),\nabla F\left(\boldsymbol{r}\right)\right)^{(\ell-1)}$, the step size uses the Barzilai–Borwein initialization
@@
$F\left(\boldsymbol{r}+\boldsymbol{\beta}\left(\alpha\right)\right)\geq F\left(\boldsymbol{r}\right)+\kappa_{1}\nabla_{\boldsymbol{r}}F\left(\boldsymbol{r}\right)^{T}\boldsymbol{\beta}\left(\alpha\right),$ (87)
with $\kappa_{1}\!\in\!(0,1)$ and the backtracking using $\alpha\!\leftarrow\!\kappa_{2}\alpha$, $\kappa_{2}\in(0,1)$. The iteration terminates when it meets the stopping criteria that the projected gradient norm is smaller than a threshold $\varepsilon_{\text{pg}}$, i.e., $\|\boldsymbol{\upsilon}_{\alpha}(\boldsymbol{r}^{(\ell)})\|_{2}\ \leq\ \varepsilon_{\text{pg}}$, or the relative step is smaller than a threshold $\varepsilon_{\text{rel}}$, i.e., $\frac{\|\boldsymbol{r}^{(\ell)}-\boldsymbol{r}^{(\ell-1)}\|_{2}}{\max\{1,\|\boldsymbol{r}^{(\ell-1)}\|_{2}\}}\ \leq\ \varepsilon_{\text{rel}}$. For box constraints, $\boldsymbol{\upsilon}_{\alpha}(\boldsymbol{r}^{(\ell)})$ admits the following component-wise form for $\imath$-th entry, according to (69), independent of $\alpha$
$\!\!\!\big[\boldsymbol{\upsilon}_{\alpha}(\boldsymbol{r}^{(\ell)})\big]_{\imath}\!=\!\!\begin{cases}[\nabla_{\boldsymbol{r}}F^{\left(\ell\right)}(\boldsymbol{r})]_{\imath},\!&\!\!\!r_{\text{lb},i}\!\!<\!r_{i}\!<\!r_{\text{ub},i},\\ \max\{0,[\nabla_{\boldsymbol{r}}F^{(\ell)}(\boldsymbol{r})]_{\imath}\},\!&\!\!\!r_{\imath}=r_{\text{lb},\imath},\\ \min\{0,[\nabla_{\boldsymbol{r}}F^{(\ell)}(\boldsymbol{r})]_{\imath}\},\!&\!\!\!r_{\imath}=r_{\text{ub},\imath}.\end{cases}\!\!\!\!$ (88)
The gradient cost is dominated by Jacobians of rotation-related steering vectors and derivatives with matrix-chain products. Because the directional-gain model is piecewise-smooth, we use subgradients in (88) along the visibility boundary to ensure stable line search.
@@
1:  Input: fixed $(\boldsymbol{W},\boldsymbol{\theta})$, box $\mathcal{B}$, tolerances $\varepsilon_{\mathrm{pg}},\varepsilon_{\mathrm{rel}}>0$, and maximum iterations $I_{\text{PGD}}^{\text{max}}$.
2:  Initialization: initial point $\boldsymbol{r}^{(0)}\in\mathcal{B}$
@@
13:   Compute projected gradient mapping $\boldsymbol{\upsilon}_{\alpha}(\boldsymbol{r}^{(\ell)})$ via (88).
14:   if $\|\boldsymbol{\upsilon}_{\alpha}(\boldsymbol{r}^{(\ell+1)})\|_{2}\leq\varepsilon_{\mathrm{pg}}$ or $\dfrac{\|\boldsymbol{r}^{(\ell+1)}-\boldsymbol{r}^{(\ell)}\|_{2}}{\max\{1,\|\boldsymbol{r}^{(\ell)}\|_{2}\}}\leq\varepsilon_{\mathrm{rel}}$ then
@@
18:  Output: $\boldsymbol{r}^{\text{B}}\leftarrow\boldsymbol{r}^{(\ell+1)}_{1:3}$, $\boldsymbol{r}^{\text{R}}\leftarrow\boldsymbol{r}^{(\ell+1)}_{4:6}$.
III-D3 Overall algorithm and complexity analysis for subproblem 3
@@
Algorithm 3 summarizes the overall projected gradient method. For the per-iteration computational complexity, directly forming $\boldsymbol{B}$ and the Jacobians as dense $M\times N$ matrices yields $\mathcal{O}(PMN)$ computations per rotation axis, and evaluating the sensing gradients over $A$ grids would add $\mathcal{O}(AMN)$. However, using the rank-$P$ factorization for $\boldsymbol{B}$ and its derivatives, all costly matrix–vector products, involving $\boldsymbol{B}$ and $\partial\boldsymbol{B}/\partial r_{\imath}^{\text{A}}$, are computed without materializing $M\times N$ blocks and costing $\mathcal{O}(NP)$ plus $\mathcal{O}(MP)$. With these reuses, one PGD iteration with full gradient of $F$ has the leading cost: i) precomputation on steering vectors and their Jacobians with $\mathcal{O}\big(P(M+N)\big)$, ii) sensing and communication channel–Jacobian products $\mathcal{O}\big((A+K)P(N+M)\big)$, iii) use $\boldsymbol{W}(\boldsymbol{W}^{H}\boldsymbol{f})$ to apply $\boldsymbol{W}\boldsymbol{W}^{H}$ in (72) and (82) with $\mathcal{O}\!\big((M+K)M(A+K)\big)$. Therefore, the per-iteration complexity scales as $\mathcal{O}\big(P(M+N)(A+K+1)+(M+K)M(A+K)\big)$ with a constant coefficient affected by backtracking searches.
III-E Overall algorithm for the original problem
Finally, Algorithm 4 summarizes the overall AO procedure for (P1). At the $\zeta$-th outer iteration, we first update the NMSE scaling parameter by $\iota^{(\zeta)}\leftarrow\iota^{\star}$ in (23). Then, with the other blocks fixed, we update $\boldsymbol{W}$ by solving (P2), update $\boldsymbol{\theta}$ by solving (P3) on the complex circle manifold, and update $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$ by solving (P4) via projected gradient ascent under box constraints. Since each block update is designed to be non-decreasing in the objective of (P1) with $\iota$ fixed, and the $\iota$-update minimizes $\text{NMSE}(\iota)$ for fixed $(\boldsymbol{W},\boldsymbol{\theta},\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$, the overall objective in (24a) is monotonically non-decreasing over outer iterations. The algorithm thus converges to a stationary point of (P1) under standard regularity assumptions. However, due to the intrinsic non-convexity of (P1), global optimality is generally intractable, and the proposed AO-based algorithm obtains a suboptimal solution for feasible system design.
Algorithm 4 AO-based Algorithm for (P1)
1:Input: channel and angular parameters, $P_{\text{B}}$, $\rho$, desired pattern $\{\mathcal{P}_{\text{d},a}\}_{a=1}^{A}$, tolerances $\epsilon_{\mathrm{AO}}$.
2:Initialize: feasible $\boldsymbol{W}^{(0)}$, $\boldsymbol{\theta}^{(0)}$, and $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})^{(0)}$ and set $\zeta=0$.
@@ III-E Overall algorithm for the original problem
4:  Compute current beampattern samples $\boldsymbol{p}^{(\zeta)}=[\mathcal{P}_{1},\ldots,\mathcal{P}_{A}]^{T}$ and update $\iota^{(\zeta)}\leftarrow\iota^{\star}$ via (23).
5:  Update $\boldsymbol{W}^{(\zeta+1)}$ by solving (P2) via QT, $M$, and bisection search.
6:  Update $\boldsymbol{\theta}^{(\zeta+1)}$ by solving (P3) via RCG on the complex circle manifold.
7:  Update $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})^{(\zeta+1)}$ by solving (P4) via projected gradient ascent with Armijo backtracking.
8:  $\zeta\leftarrow\zeta+1$.
9:until the relative increase of the objective in (P1) is below $\epsilon_{\mathrm{AO}}$ (or a maximum iteration is reached).
10:Output: $(\boldsymbol{W},\boldsymbol{\theta},\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$.
@@ IV Simulation Results
This section evaluates the proposed low-altitude ISAC design with cooperatively rotatable BS and RIS arrays. For performance characterization, we report three metrics: the downlink sum rate $R\triangleq\sum_{k=1}^{K}\log_{2}(1+\mathrm{SINR}_{k})$, the beampattern mismatch quantified by the NMSE in (22) with $\iota$ updated by (23) at each AO outer iteration, and the weighted ISAC utility $\mathcal{U}\triangleq\sum_{k=1}^{K}\log_{2}(1+\mathrm{SINR}_{k})-\rho\text{NMSE}$ of (P1).
IV-A Simulation Setup and Benchmarks
Unless otherwise stated, we set $\rho=10$. The BS located at (0,0,0)(0,0,0)employs a $M=M_{\text{col}}M_{\text{row}}=2\times 2$ UPA with inter-antenna spacing $\lambda/2$ at carrier wavelength $\lambda=0.1$ m. The RIS is a $6\times 6$ UPA with element spacing $\lambda/2$, located at $(10,0,0)$ m relative to the BS. The small-scale fading of all links follows a rotation-aware geometry-based multipath model: each BS-user and RIS-user channel contains $L_{k}=Q_{k}=2$ paths, and the BS-RIS channel contains $P=2$ paths. The azimuth and elevation angles are independently generated with $\varphi\in[-\pi,\pi]$ and $\vartheta\in[-\pi/3,\pi/3]$, and the complex path gains are drawn from independent circularly symmetric Gaussian distributions with normalized variance. For sensing, we discretize the angular region $\phi\in[-\pi,\pi]$ and $\vartheta\in[-\pi/4,\pi/4]$ with $A_{\mathrm{az}}=11$ azimuth sampling points and $A_{\mathrm{el}}=6$ elevation points, resulting in $A=A_{\mathrm{az}}A_{\mathrm{el}}=66$ grid points. The desired pattern is a spotlight-type template consisting of two angular sectors, and the NMSE is computed with the optimal scaling $\iota^{\star}$ in (23). The maximum rotation angle of the active and passive array is set by $\pi/2$ for each Euler angle in (24d) and (24e) by $r_{\text{i},\min}^{\text{R}}=r_{\text{i},\min}^{\text{B}}=-\frac{\pi}{2},r_{\text{i},\max}^{\text{R}}=r_{\text{i},\max}^{\text{B}}=\frac{\pi}{2},\text{i}\in\{\mathrm{x},\mathrm{y},\mathrm{z}\}$. We consider two settings on the directional radiation pattern gain of the BS antenna and the RIS element: half-space isotropic ($b^{\text{B}}=b^{\text{R}}=0$) and directional ($b^{\text{B}}=b^{\text{R}}=2$), where the latter yields a narrower mainlobe and stronger gain variations over directions.
We compare the proposed jointly rotatable BS and RIS scheme (“Rot-BS & Rot-RIS”) with the following baselines: (i) “Rot-BS & Fix-RIS”, where only the BS array is rotatable while the RIS array orientation is fixed, (ii) “Fix-BS & Rot-RIS”, where only the RIS array is rotatable while the BS array orientation is fixed, (iii) “Fix-BS & Fix-RIS”, where both BS and RIS arrays are fixed, (iv) “Rot-BS & No-RIS”, where the RIS–user link is disabled by setting the RIS–user path gains to zero, and (v) “Fix-BS & No-RIS”. All reported curves are obtained by averaging the communication and sensing metrics over 100 independent channel realizations.